# MetaEarth: A Generative Foundation Model for Global-scale Remote Sensing Image Generation

### v2

@@ 1 Introduction
The ongoing advancement of generative foundation models has propelled natural image generation techniques to a new level. These encompass text-to-image generation [1, 2, 3], video generation [4, 5], image translation [6, 7, 8], image editing [9, 10], etc, contributing to many practical applications, including art design [11], training image augmentation [12, 13, 14], and building generative world models [15, 16, 17].
Figure 1: We propose MetaEarth, a generative foundation model that simulates Earth’s visuals from an overhead perspective. MetaEarth shows powerful capabilities on generating worldwide, multi-resolution, unbounded, and virtually limitless remote sensing images. For more visualization and animated results, please refer to our project page.
Despite the recent progress in image generation quality and diversity, the scale of the scene generated is still constrained to daily human activity scenarios, resulting in a single generated image with limited resolution and limited information capacity. In this paper, we aim to push the boundaries of generative models by extending image generation capabilities from localized human daily scenes to a global scale. We employ an overhead remote sensing perspective as a novel viewpoint to study this problem and present MetaEarth - a foundational generative model capable of generating worldwide, multi-resolution, unbounded, and virtually limitless remote sensing images.
@@ 2.3 Remote Sensing Image Generation
In the field of remote sensing, image generation models have recently garnered increasing attention. Previous research on this topic mainly focuses on text-to-image and image-to-image tasks. For the text-to-image generation task, one common approach is to fine-tune latent diffusion models [1] directly based on remote sensing text-image datasets [32, 33, 34]. The image-text data pairs are usually obtained by a pre-trained image captioning model and then are refined manually assisted by GPT-4. In addition to the direct fintuning, Sebaq et al. [35] propose a generation framework containing two individual diffusion models, where the first model generates images from prompts and the second one upscales the generated images. For the image-to-image generation task, the input conditional images usually include maps [36] and semantic layouts [37, 38, 39].
@@ 2.4 Remote Sensing Foundation Model
Remote sensing vision-language foundation models are trained on a large amount of image-text paired data. Some methods [49, 50, 51] adopt contrastive learning, which aims to train an image encoder and a text encoder to align the visual and linguistic domains. The learned aligned vision-language representations can be used for many downstream tasks, such as zero-shot classification [49], image-text retrieval [52, 53], and image captioning [54, 55]. For example, Zhang et al. [50] filtered remote sensing image-text pairs from publicly available large datasets to build the RS5M dataset and fine-tuned the CLIP [56] model to obtain a GeoRSCLIP model. Recently, autoregressive pre-training has been widely used for training vision-language models, such as [57, 58, 59, 60]. Most methods utilize CLIP to extract visual features and design connectors to map these visual features into large language models for predicting corresponding text. They typically start with image-text pre-training and then further refine the model through instruction fine-tuning.
@@ 3.3 Unbounded Image Generation
To achieve the generation of large-scale remote sensing images of arbitrary sizes, we propose an unbounded image generation method including a memory-efficient sliding window generation pipeline and a noise sampling strategy. Figure 3 illustrates the proposed method.
@@ 3.4.2 Training Details
where ${k}$, ${r}$, ${n}$, ${[\cdot]_{JPEG}}$ represent the blur kernel, the scaling factor, noise, JPEG compression, respectively. During training, we utilize high-order degradation to downsample real images into corresponding low-resolution ones, following the parameter settings as described in [63]. During inference, we do not apply any degradation - the low-resolution conditional input can be either real images or images generated from the previous cascading steps.
@@
where ${k}$ and ${\gamma}$ are hyperparameters, $\text{SNR}(x_{t})$ is the signal-to-noise ratio of the noisy image $x_{t}$.
Our model is implemented using PyTorch. We utilize the AdamW optimizer with an initial learning rate of 2e-6 to optimize parameters and set the total number of epochs to 30. To increase the training batch size, we employ gradient accumulation and multi-GPU parallelism. The batch size per GPU is set to 1, and gradient accumulation is set to 8. The training process takes in a total of 2,000 GPU hours on a set of NVIDIA RTX4090 GPUs. The training is initialized from scratch without loading pre-trained parameters.
@@ 4.1.1 Dataset
To generate multi-resolution remote sensing images for any region globally, we collected a large dataset, consisting of three levels of remote sensing images from Google Earth, corresponding to resolutions of 64m/pixel, 16m/pixel, and 4m/pixel, covering the vast majority of the Earth’s landmass. A total of about 731,000 non-overlapping images, each sized 256×256 pixels, were used to train and evaluate the model. For each image in the dataset, the spatial resolution and the latitude and longitude are recorded. We partition the dataset into training, validation, and testing sets in an 8:1:1 ratio.
@@ 4.2.3 Arbitrary-sized Image Generation
Figure 8: Arbitrary-sized images generated by our method.
@@
TABLE I: Ablation studies on the spatial resolution guidance. We divided the Earth into different regions and assessed the quality (FID, a lower score indicates better) of images generated in each region. ${lng}$ denotes the longitude.
@@
Fig. 8 shows the unbounded generation results with the proposed sliding window generation. Although we applied block-by-block generation, it can be observed that the proposed method generates continuous land features across image blocks. This indicates that our window overlapping and noise sampling strategy greatly alleviate visual discontinuities caused by image block stitching, thereby achieving boundless and arbitrary-sized image generation. Fig. 9 shows an example of a high-resolution large-scale image generated by our model.
@@ 4.2.4 Generalization to Unseen Conditions
Thanks to being trained on large-scale data, our model possesses strong generalization capabilities and performs well even on unseen scenes. To verify this, we designed an interesting experiment. In the experiment, we aimed to create a “parallel world” and used a low-resolution map of “Pandora Planet” (generated by GPT4-V) as the initial condition for our model and then generated higher-resolution images sequentially. Fig. 10 shows the generated results. It can be observed that despite our training data not covering such scenes, MetaEarth is still able to generate images with reasonable land cover distribution and realistic details.
@@
Figure 11: Comparision between MetaEarth to other text-to-image models including GPT-4 V, the latest version of Stable Diffusion, and Ernie. For the three compared methods, the text prompts are given as: “Please generate a 4m/pixel-resolution satellite remote sensing image of an urban scene. The image includes detailed city infrastructure such as roads, buildings, parks, and waterways with clear visibility of the urban layout and structure.”
@@ 4.2.5 Comparison with other text-to-image models
Despite the rapid advancements in generative models represented by Stable Diffusion and DALLE, their capabilities in generating overhead remote sensing scenes are limited. In Fig. 11, we compare MetaEarth with other state-of-the-art text-to-image models, including GPT-4 V (DALLE), the latest version of Stable Diffusion, and Ernie. We can see that MetaEarth significantly outperforms these models in terms of image realism and the rationality of scene layouts. The primary reason behind this difference is the lack of relevant remote sensing training data during the training phase of previous models. In addition, the three contrasting models lack sensitivity to resolution control and are difficult to accurately capture the correspondence between resolution and image content, making it difficult to generate multi-resolution images or unbounded scenes. MetaEarth, on the other hand, excels in these aspects, demonstrating its superior capability in generating high-quality, realistic remote sensing images across various resolutions and scene layouts.
@@ 4.3 Ablation Study
In this section, we conduct experiments to validate the effectiveness of the proposed methods, including the resolution-guided generation and unbounded generation.
4.3.1 Effectiveness of Resolution Guidance
@@
We conducted quantitative analysis in different regions around the world. The regions are divided by using different criteria, including the division of regions at different longitudes, or the division of northern and southern hemispheres. The experimental results are shown in Table I. The “w/o sr” rows represent the network without using the resolution guidance. The first column of the table gives the resolution of the generated images. Except for the 4m*, the model inputs are low-resolution images obtained by downsampling the corresponding resolution real images through bicubic interpolation. The inputs for 4m* are randomly cropped from the real 16m resolution images. Due to the different characteristics of sensors of different resolution levels, the FID value of 4m* is significantly higher than that of other rows, but our resolution guidance strategy is still effective.
@@ 4.3.2 Evaluation on Noise Sampling
To measure the continuity of the different generated blocks in the seam area, in Fig. 3, we compare the generated results w/ and w/o using the proposed noise sampling strategy. We can see that with the proposed noise sampling strategy, we obtained smoother results, and the traces of patchwork are difficult to detect with the naked eye. We also designed a gradient-based method to quantitatively evaluation the transition smoothness of different approaches. We randomly sample 750 generated images and calculate the absolute directional gradient values of adjacent pixels at the seams and then take the average. Experimental results are shown in Table II. We evaluate the proposed two sub-strategies respectively, one is overlapped sliding window (marked as “Overlap” in Table II) and another is the constraints on initial noise sampling (marked as “Noise Constraint” in Table II). Smaller gradient values indicate better results. The experimental results demonstrate that the combined use of the overlapping cropping strategy and our proposed noise sampling method can effectively alleviate the seams appearing during image block stitching, thereby enhancing the visual quality of the generated arbitrary-sized images.
TABLE II: Ablation experiment on our proposed noise sampling method. We compare the averaged directional gradient value of pixels in the window boundary area (a lower value indicates better).
@@ 4.4 Evaluation on Downstream Tasks
The proposed MetaEarth can be used as a powerful data engine to generate high-quality training data that can benefit downstream tasks. In this experiment, we choose remote sensing image classification, a fundamental task in remote sensing image interpretation, as our test bed for downstream evaluation.
To quantitatively demonstrate the application of our model in downstream tasks and further evaluate the fidelity and diversity of the generated images, we design a remote sensing image classification experiment, using our model as a data engine to provide training data support. Our classification dataset contains seven categories of remote sensing scenes: beach, desert, farmland, forest, industrial area, mountain, and residential area. Each class contains approximately 150 images, with image sizes of 256$\times$256 pixels. The dataset is re-collected from our global 4m resolution images. The training and testing sets are divided by a 3:1 ratio.
We choose four different models as classifiers: VGG19 [67], ResNet34 [68], ViT-B/32, and ViT-B/16 [69]. We first train four baseline models using the original dataset. Then, we use the generation model to augment the training dataset to five times its original size. After mixing the augmented data with the original one, we retrain the four models from scratch using the combined dataset. For the augmentation process, we downsample the images in the training dataset and utilize the model to generate diverse results.
TABLE III: Accuracy of the downstream remote sensing image classification task w/ and w/o image augmentation.
@@
Table III shows the classification accuracy on the test set for different models using different training data. Compared with the original models, the models trained with data augmentation show consistently improved classification accuracy. This indicates that MetaEarth provides informative data for training, which also reflects the high quality and diversity of the images generated. This indicates that when faced with tasks with insufficient or sparse data, MetaEarth has the potential to serve as a data engine to provide usable data at lower costs.
4.5 Disccusion

### v3

@@ 1 Introduction
The ongoing advancement of generative foundation models has propelled natural image generation techniques to a new level[1]. These encompass text-to-image generation [2, 3, 4], video generation [5, 6], image translation [7, 8, 9], image editing [10, 11], etc, contributing to many practical applications, including art design [12], training image augmentation [13, 14, 15], and building generative world models [16, 17, 18].
Figure 1: We propose MetaEarth, a generative foundation model that simulates Earth’s visuals from an overhead perspective. MetaEarth shows powerful capabilities on generating worldwide, multi-resolution, unbounded, and virtually limitless remote sensing images. For more visualization and animated results, please refer to our project page: https://jiupinjia.github.io/metaearth/.
Despite the recent progress in image generation quality and diversity, the scale of the scene generated is still constrained to daily human activity scenarios, resulting in a single generated image with limited resolution and limited information capacity. In this paper, we aim to push the boundaries of generative models by extending image generation capabilities from localized human daily scenes to a global scale. We employ an overhead remote sensing perspective as a novel viewpoint to study this problem and present a foundational generative model MetaEarth. Our methods innovatively propose a self-cascading framework and a noise sampling strategy, enabling MetaEarth to generate worldwide, multi-resolution, unbounded, and virtually limitless remote sensing images. Fig. 1 showcases some results of our model.
@@ 2.3 Remote Sensing Image Generation
In the field of remote sensing, image generation models have recently garnered increasing attention. As shown in Table I, previous research on this topic mainly focuses on text-to-image and image-to-image tasks.
TABLE I: Comparison of existing generative models for remote sensing images.
Method Control Signal Global-scale Resolution Range Image Size (pixels)
DiffusionSat [35] text, metadata $\times$ limited 512$\times$512
Crs-diff [36] text, metadata image $\times$ limited 224$\times$224
RSFSG-Diff [37] semantic map $\times$ fixed 256$\times$256
MapSat [38] map $\times$ fixed 256$\times$256
MetaEarth (ours) image, metadata $\checkmark$ multi-resolution unbounded
For the text-to-image generation task, one common approach is to fine-tune latent diffusion models [4] directly based on remote sensing text-image datasets [39, 35, 36]. The image-text data pairs are usually obtained by a pre-trained image captioning model and then are refined manually assisted by GPT-4. In addition to the direct fintuning, Sebaq et al. [40] propose a generation framework containing two individual diffusion models, where the first model generates images from prompts and the second one upscales the generated images.
The image-to-image generation task is primarily divided into two categories: layout-to-image [41] and modality transfer [42, 43]. The former’s input conditional images typically include maps [38] and semantic layouts [37, 44], while the latter aims to achieve the transformation between multimodal data such as RGB images, infrared images, SAR images, and hyperspectral images.
@@ 2.4 Remote Sensing Foundation Model
Remote sensing vision-language foundation models are trained on a large amount of image-text paired data. Some methods [54, 55, 56] adopt contrastive learning, which aims to train an image encoder and a text encoder to align the visual and linguistic domains. The learned aligned vision-language representations can be used for many downstream tasks, such as zero-shot classification [54], image-text retrieval [57, 58], image captioning [59, 60], and change captioning [61, 62]. For example, Zhang et al. [55] filtered remote sensing image-text pairs from publicly available large datasets to build the RS5M dataset and fine-tuned the CLIP [63] model to obtain a GeoRSCLIP model. Recently, autoregressive pre-training has been widely used for training vision-language models, such as [64, 65, 66, 67]. Most methods utilize CLIP to extract visual features and design connectors to map these visual features into large language models for predicting corresponding text. They typically start with image-text pre-training and then further refine the model through instruction fine-tuning.
@@ 3.3 Unbounded Image Generation
To achieve the generation of large-scale remote sensing images of arbitrary sizes, we propose an unbounded image generation method including a memory-efficient sliding window generation pipeline and a noise sampling strategy. Fig. 3 illustrates the proposed method.
@@ 3.4.2 Training Details
where ${k}$, ${r}$, ${n}$, ${[\cdot]_{JPEG}}$ represent the blur kernel, the scaling factor, noise, JPEG compression, respectively. During training, We directly apply the high-order degradation, composed of two simple degradation models, in the pixel domain of high-resolution images to downsample real images into corresponding low-resolution ones, i.e. $n=2$:
$x_{lr}=\mathcal{D}^{2}(x_{hr})=({\mathcal{D}_{1}}\circ{\mathcal{D}_{2}})(x_{hr})$ (25)
Following the parameter settings as described in [70], in terms of blur kernels, we chose isotropic and anisotropic Gaussian blur kernels, as well as sinc filters. During the scaling phase, we randomly used nearest, bilinear, and bicubic interpolation methods to enlarge or downscale the images; for noise, we randomly added Gaussian or Poisson noise to the images, with the noise type randomly selected as colour or grey; for the JPEG compression process, we randomly compressed the images with quality factors ranging from 30 to 95.
During inference, we do not apply any degradation - the low-resolution conditional input can be either real images or images generated from the previous cascading steps.
@@
In this equation, $k$ and $\gamma$ are hyperparameters that were both set to 1 in the experiments, following the suggestions in  [71]. The term $\text{SNR}(x_{t})$ represents the signal-to-noise ratio of the noisy image.
Our model is implemented using PyTorch. We utilize the AdamW optimizer with an initial learning rate of 2e-6 to optimize parameters and set the total number of epochs to 30. To increase the training batch size, we employ gradient accumulation and multi-GPU parallelism. The batch size per GPU is set to 1, and gradient accumulation is set to 8. The training process takes more than 2,000 GPU hours on a set of NVIDIA RTX4090 GPUs. The training is initialized from scratch without loading pre-trained parameters.
@@ 4.1.1 Dataset
To generate multi-resolution remote sensing images for any region globally, we collected a large dataset, which includes three levels of 256$\times$256 sized remote sensing images at arbitrary latitudes and longitudes from Google Earth, corresponding to resolutions of 64m/pix, 16m/pix, and 4m/pix. These data encompass nearly all of the Earth’s geographical and environmental conditions, including cities, forests, deserts, oceans, glaciers, and more. Then, we cleansed the dataset, removing some ocean images that were highly repetitive, as there was no need to use an excessive amount of similar data for model training. We also eliminated images that contained noise or significant cloud cover. Lastly, for the handpicked dataset, we randomly sampled approximately 1,000,000 non-overlapping images at each resolution (a total of about 3,100,000 images) for model training, and additionally randomly sampled about 140,000 images, which were divided into a validation set and a test set in a 1:1 ratio. The training set images at various resolutions are shown in Table II.
TABLE II: The number of training data at various resolutions.
Spatial Resolution Image Count
64m/pixel 916,509
16m/pixel 1,022,194
4m/pixel 1,186,312
We strictly adhered to Google Earth’s copyright and guidelines during our data collection process, and upon acceptance of the paper, we deleted all the training data.
@@ 4.2.3 Arbitrary-sized Image Generation
TABLE III: Ablation studies on the spatial resolution guidance. We divided the Earth into different regions and assessed the quality (FID, a lower score indicates better) of images generated in each region. ${lng}$ denotes the longitude.
@@
Fig. 8 shows some examples of high-resolution large-scale images generated with the proposed sliding window generation. Although we applied block-by-block generation, it can be observed that the proposed method generates continuous land features across image blocks. This indicates that our window overlapping and noise sampling strategy greatly alleviate visual discontinuities caused by image block stitching, thereby achieving boundless and arbitrary-sized image generation.
@@ 4.2.4 Generalization to Unseen Conditions
Figure 10: The self-cascaded generation results from MetaEarth with image sourced from Sentinel-2 satellite.
Thanks to being trained on large-scale data, our model possesses strong generalization capabilities and performs well even on unseen scenes. We conducted two sets of experiments to verify the generalization capability of our MetaEarth model on unseen conditions. In the first experiment, we aimed to create a “parallel world” and used a low-resolution map of “Pandora Planet” (generated by GPT-4V) as the initial condition for our model and then generated higher-resolution images sequentially. As shown in Fig. 9, MetaEarth is capable of generating images that have a reasonable land cover distribution along with realistic details. In the second experiment, we utilized real remote sensing images collected by the Sentinel-2 satellite as low-resolution inputs to generate multi-resolution images, with the results depicted in Fig. 10. The generated images preserved the imaging style of the original sensor while exhibiting distinct regional characteristics at various resolutions. It is noted that despite our training data not covering scenes in these two experiments, MetaEarth is still able to generate clear, semantically reasonable, and regionally styled realistic images, which indicates that our model possesses strong generation capabilities and performs well even on unseen scenes.
@@ 4.2.5 Comparison with other text-to-image models
Despite the rapid advancements in generative models represented by Stable Diffusion and DALLE, their capabilities in generating overhead remote sensing scenes are limited. In Fig. 11, we compare MetaEarth with other state-of-the-art text-to-image models, including GPT-4V (DALLE), the latest version of Stable Diffusion, and Ernie. We can see that MetaEarth significantly outperforms these models in terms of image realism and the rationality of scene layouts. The primary reason behind this difference is the lack of relevant remote sensing training data during the training phase of previous models. In addition, the three comparison models lack sensitivity to resolution control and are difficult to accurately capture the correspondence between resolution and image content, making it difficult to generate multi-resolution images or unbounded scenes. MetaEarth, instead, excels in these aspects, demonstrating its superior capability in generating high-quality, realistic remote sensing images across various resolutions and scene layouts.
Figure 11: Comparision between MetaEarth to other text-to-image models including GPT-4V (a), the latest version of Stable Diffusion (b), and Ernie (c). For the three compared methods, the text prompts are given as: “Please generate a 4m/pixel-resolution satellite remote sensing image of an urban scene. The image includes detailed city infrastructure such as roads, buildings, parks, and waterways with clear visibility of the urban layout and structure.”
@@ 4.3 Ablation Study
In this section, we conduct experiments to validate the effectiveness of the proposed methods, including the self-cascaded generation framework, resolution-guided generation and unbounded generation.
4.3.1 Necessity of Self-cascaded Generation Framework
The self-cascaded generation framework we propose aims to generate multi-resolution images in a stage-by-stage manner. In addition, we have explored an alternative one that is fundamentally different from our self-cascaded method, which involves generating the highest-resolution image by splicing directly and then downsampling it level-by-level to get each low-resolution image. This method contrasts with the self-cascaded generation method presented in this paper, representing two distinct framework for image generation.
To verify the necessity and rationality of self-cascaded generation framework within the limits of maximum hardware feasibility, we developed a generative model that is structurally and parametrically close to MetaEarth, with an input size of 4$\times$4 and an output size of 256$\times$256, capable of directly generating 4m resolution images from 256m resolution imagery. Employing our simplified unbounded image generation method, the large-sized image with 4m resolution generated directly by this model and the one generated using the MetaEarth cascaded approach are shown in Fig. 12.
Figure 12: Comparison between directly generating the highest resolution image (a) and generating it in a stage-by-stage manner (b). The generated images have a resolution of 4m/pix and a size of 1024$\times$1024 pixels. With the red-boxed region from the first column of images as input, it is evident that the cascaded generation image exhibits more distinct semantics, clearer features, and a more rational layout than the directly generated one.
It is evident that the image generated through a cascaded approach clearly surpasses the directly generated one in visual quality, exhibiting more distinct semantics, sharper feature contours, and a more rational layout. The structural and global semantic information provided by the 4$\times$4-sized images is severely lacking, which will lead to two aspects of problems. On the one hand, this makes it difficult for the model to produce high-quality images with clear features and a sensible distribution of features. On the other hand, this allows the model to be approximated as an unconditional generative model, where the generation results are primarily determined by the initial noise, with the conditional image only having a minor impact. With our simplified unbounded image generation method, the identical initial noise leads to similar outputs across image blocks, resulting in the failure of the algorithm to compile them into a semantically coherent, continuous large-scale image. Considering these points, we therefore conclude that directly generating the highest resolution images is not as sensible an approach, and that employing a cascaded generation framework is essential.
4.3.2 Effectivene of Resolution Guidance
@@
We conducted quantitative analysis in different regions around the world. The regions are divided by using different criteria, including the division of regions at different longitudes, or the division of northern and southern hemispheres. The experimental results are shown in Table III. The “w/o sr” rows represent the network without using the resolution guidance. The first column of the table gives the resolution of the generated images. Except for the 4m*, the model inputs are low-resolution images obtained by downsampling the corresponding resolution real images through bicubic interpolation. The inputs for 4m* are randomly cropped from the real 16m resolution images. Due to the different characteristics of sensors of different resolution levels, the FID value of 4m* is significantly higher than that of other rows, but our resolution guidance strategy is still effective.
@@ 4.3.3 Evaluation on Noise Sampling
To measure the continuity of the different generated blocks in the seam area, in Fig. 3, we compare the generated results w/ and w/o using the proposed noise sampling strategy. We can see that with the proposed noise sampling strategy, we obtained smoother results, and the traces of patchwork are difficult to detect with the naked eye. We also designed a gradient-based method to quantitatively evaluation the transition smoothness of different approaches. We randomly sample 750 generated images and calculate the absolute directional gradient values of adjacent pixels at the seams and then take the average. Experimental results are shown in Table IV. We evaluate the proposed two sub-strategies respectively, one is overlapped sliding window (marked as “Overlap” in Table IV) and another is the constraints on initial noise sampling (marked as “Noise Constraint” in Table IV). Smaller gradient values indicate better results. The experimental results demonstrate that the combined use of the overlapping cropping strategy and our proposed noise sampling method can effectively alleviate the seams appearing during image block stitching, thereby enhancing the visual quality of the generated arbitrary-sized images.
TABLE IV: Ablation experiment on our proposed noise sampling method. We compare the averaged directional gradient value of pixels in the window boundary area (a lower value indicates better).
@@
4.3.4 Necessity of High-order Degradation Operation
Figure 14: A comparison between the generated images w/ and w/o using the high-order degradation. Both models take real images with a spatial resolution of 16m/pix as input.
In the training process, we introduce high-order degradation operation to address the challenges of shifting between different distributions. In this section, we have designed a new ablation experiment to validate the necessity of the high-order degradation operation, where we compared the simple downsampling method (bicubic interpolation) with the high-order degradation strategy.
The experiment primarily investigates the single-stage generation process from 16m to 4m resolution, using both bicubic interpolation and high-order degradation to obtain low-resolution images from high-resolution ones for training. Under the same model architecture, parameter volume, training data, and other experimental conditions, we examine the performance of the two generated models in response to different inputs.
As shown in Fig. 14, when both models process real remote sensing images, the one trained with high-order degradation method generates clearer imagery, whereas the model relying on simple downsampling method results in significantly blurred images. For a well-trained model, the distribution of its output images is similar to that of the real data, which means that the model without high-order degradation cannot take the images generated by the previous stage of the model as input to cascade and generate higher-resolution images, and thus cannot adapt to the self-cascaded framework.
We also calculated the FID for the images generated by the two models. It can be observed from the Table V that our method exhibits a notable advantage in terms of FID, signifying that the high-order degradation we implemented effectively boosts the model’s ability to generalize to real input data, resulting in the generation of superior quality high-resolution images.
Both qualitative and quantitative analyses demonstrate that the high-order degradation operation we have adopted is necessary, enhancing the model’s adaptability to real data inputs and enabling the model to be applied in a self-cascaded generative framework.
TABLE V: Ablation experiment on high-order degradation operation.
[…]
$\times$ 148.31
${\bm{{78.72}}}$
@@ 4.4 Evaluation on Downstream Tasks
The proposed MetaEarth can be used as a powerful data engine to generate high-quality training data that can benefit downstream tasks. For instance, MetaEarth enables data augmentation by employing a process of downsampling followed by upsampling, providing a vast amount of diverse data to enhance classification tasks that suffer from insufficient or sparse data. In addition, MetaEarth can be utilized to synthesize hybrid scenes that combine real objects and virtual background for object detection, especially for tasks involving rare or scarce targets. Furthermore, MetaEarth may have great potential to build deepfake detection algorithms in the remote sensing field, maintaining the security of geographic information systems and preventing false data from misleading decision-making. With MetaEarth’s robust generation capabilities across a global scale and various resolutions, users can obtain multi-scale images with different regional characteristics. This may further enhance the generalization of deepfake detection models and improve their ability to distinguish between real and fake data.
In this paper, to quantitatively demonstrate the application of our model in downstream tasks and further evaluate the fidelity and diversity of the generated images, we choose remote sensing image classification, a fundamental task in remote sensing image interpretation, as our test bed for downstream evaluation. In this experiment, we use our model as a data engine to provide training data support. Our classification dataset contains seven categories of remote sensing scenes: beach, desert, farmland, forest, industrial area, mountain, and residential area. Each class contains approximately 150 images, with image sizes of 256$\times$256 pixels. The dataset is re-collected from our global 4m/pix resolution images. The training and testing sets are divided by a 3:1 ratio.
Figure 15: The generated augmented data. The leftmost column displays input conditional images, with the top-to-bottom rows representing industrial areas, farmlands, and residential areas, respectively. The right three columns show images generated from the real images.
We choose four different models as classifiers: VGG19 [74], ResNet34 [75], ViT-B/32, and ViT-B/16 [76]. We first train four baseline models using the original dataset. Then, we use the generation model to augment the training dataset to five times its original size. After mixing the augmented data with the original one, we retrain the four models from scratch using the combined dataset. For the augmentation process, we downsample the images in the training dataset and utilize the model to generate diverse results. Some of the augmented results are shown in Fig 15.
TABLE VI: Accuracy of the downstream remote sensing image classification task w/ and w/o image augmentation.
@@
Table VI shows the classification accuracy on the test set for different models using different training data. Compared with the original models, the models trained with data augmentation show consistently improved classification accuracy. This indicates that MetaEarth provides informative data for training, which also reflects the high quality and diversity of the images generated. This indicates that when faced with tasks with insufficient or sparse data, MetaEarth has the potential to serve as a data engine to provide usable data at lower costs.
4.5 Limitation
Figure 16: Failure cases of MetaEarth in generating small-sized objects. The first and second rows show unsuccessful results in generating airplanes and ships.
Currently MetaEarth may not be able to generate small objects such as aircraft and ships well. Some failure cases are shown in Fig. 16, where the shapes of the planes or ships are almost completely lost, and their features are hard to discern. This is due to the fact that, constrained by computational resources, our highest resolution for training data is 4m/pix. At this resolution, the targets are so small and occupy very few pixels, leading to a significant issue of data imbalance. During training, background patterns become the dominant focus of learning, making it challenging for the model to accurately capture the characteristics of the targets. In future research, to address these issues, we will continue to explore the potential of remote sensing generative models to achieve the generation of higher resolution images and controllable embedding of targets. We will also attempt to expand the generated modalities, not limited to the visible light images, and further investigate the controllable generation of multimodal data such as multispectral and DEM images.
5 Discussion